# Uniqueness of Galilean and Carrollian limits of gravitational theories and application to higher derivative gravity

### v3

@@ I Introduction
While calculating the Carrollian limit of Maxwell’s theory, it was shown that there are two nonequivalent Carrollian limits for any Lorentzian theory. The first one keeps the electric part in Maxwell’s theory and so was named the “electric limit”, and the other keeps the magnetic part and so was called the “magnetic limit”. There are multiple ways to take the (electric and magnetic) Carrollian limit of a Lorentzian gravitational theory:
@@
In this paper, we show that the PNR parametrization can be constructed from GAG. Similarly, the PUL parametrization can be constructed from CAG. We also show that the PUL and ZS approaches of taking the Carrollian limit are completely equivalent (and, similarly, PNR and IS). Throughout this paper, we consider the ADM decomposition with an extra real parameter $\epsilon$ which can be interpreted as a signature if it is $1$, $-1$, or 00; the zero signature case is a Carrollian manifold [79]. Once we establish the uniqueness of the Galilean and Carrollian limits (i.e., the equivalence of above approaches), we construct an algorithm for computing the $n$-th order of the Galilean, and the Carrollian expansions of a completely generic higher derivative gravity (HDG). The problem is transformed into computationally easier optimization problems which upon solving gives the desired order.
@@ II.4 Infinite scaling approach
Starting from a Lorentzian theory and writing the metric in the same form as in the ZS approach but taking the limit of the norm of the normal vector to $\infty$- ∞ results in a Galilean theory. Alternatively, starting with the Riemannian theory and taking the limit to $\infty$ results in the same Galilean theory; the difference would be only in an overall sign in the Lagrangian.
@@ II.5 Carroll algebra gauging
$H\to\tau_{\mu},\hskip 10.0pt$ $P_{a}\to e_{\mu}{}^{a},$ $G_{a}\to\omega_{\mu}{}^{a},$ $\hskip 10.0ptJ_{ab}\to\omega_{\mu}{}^{ab}.$ (11)
@@
$\zeta=\xi^{\mu}\tau_{\mu},\hskip 10.0pt$ $\zeta^{a}=\xi^{\mu}e_{\mu}{}^{a},$ $\hskip 10.0pt\lambda^{a},\hskip 5.0pt$ $\hskip 10.0pt\lambda^{ab}.$ (13)
@@
$F_{\mu\nu}=\partial_{\mu}\mathcal{A}_{\nu}-\partial_{\nu}\mathcal{A}_{\mu}+[\mathcal{A}_{\mu},\mathcal{A}_{\nu}]=HR_{\mu\nu}(H)+P_{a}R_{\mu\nu}{}^{a}(P)+G_{a}R_{\mu\nu}{}^{a}(G)+\tfrac{1}{2}J_{ab}R_{\mu\nu}{}^{ab}(J),$ (15)
@@
$R_{\mu\nu}(H)$ $=2\partial_{[\mu}\tau_{\nu]}-2\omega_{[\mu}{}^{a}e_{\nu]a},$ $\hskip 10.0ptR_{\mu\nu}{}^{a}(P)$ $=2\partial_{[\mu}e_{\nu]}{}^{a}-2\omega_{[\mu}{}^{ab}\omega_{\nu]b},$ (16)
$R_{\mu\nu}{}^{a}(G)$ $=2\partial_{[\mu}\omega_{\nu]}{}^{a}-2\omega_{[\mu}{}^{ab}\omega_{\nu]b},$ $\hskip 10.0ptR_{\mu\nu}{}^{ab}(J)$ $=2\partial_{[\mu}\omega_{\nu]}{}^{ab}-2\omega_{[\mu}{}^{a}_{c}\omega_{\nu]}{}^{cb}.$
@@ II.6 Galilei algebra gauging
$R_{\mu\nu}(H)$ $=2\partial_{[\mu}\tau_{\nu]},$ $R_{\mu\nu}{}^{a}(P)$ $=2\partial_{[\mu}e_{\nu]}{}^{a}-2\omega_{[\mu}{}^{ab}e_{\nu]b}-2\omega_{[\mu}{}^{a}\tau_{\nu]},$ (18)
$R_{\mu\nu}{}^{a}(G)$ $=2\partial_{[\mu}\omega_{\nu]}{}^{a}-2\omega_{[\mu}{}^{ab}\omega_{\nu]b},$ $\hskip 10.0ptR_{\mu\nu}{}^{ab}(J)$ $=2\partial_{[\mu}\omega_{\nu]}{}^{ab}-2\omega_{[\mu}{}^{ac}\omega_{\nu]}{}^{b}_{c},$
@@ III.1 Equivalence of PUL and ZS approaches
[…]
@@
where $A_{1}^{\mu}=-n^{\mu}\partial_{\nu}n^{\nu}+n^{\nu}\partial_{\nu}n^{\mu}-\tfrac{1}{2}n^{\mu}n^{\sigma}h^{\alpha\rho}\big{(}2\partial_{(\alpha}h_{\sigma)\rho}-\partial_{\rho}h_{\alpha\sigma}\big{)}+h^{\mu\rho}n^{\sigma}n^{\nu}\partial_{\nu}h_{\sigma\rho}-\tfrac{1}{2}n^{\sigma}n^{\mu}h^{\mu\rho}\partial_{\rho}h_{\sigma\nu}$, and $A_{2}^{\mu}=-h^{\mu\rho}n^{\sigma}\partial_{\sigma}n_{\rho}$, where $\partial_{\mu}$ is the partial derivative with respect to a generic coordinate system. This formula is the same as the PUL parameterization of the Ricci scalar up to change of connections and renaming quantities.
@@ III.3 Equivalence of PUL and CAG approaches
$v^{\mu}\tau_{\mu}=-1,$ $\hskip 10.0ptv^{\mu}e_{\mu}{}^{a}=\tau_{\mu}e^{\mu}{}_{b}=0,\hskip 10.0pt$ $e_{\mu}{}^{a}e^{\mu}{}_{b}=\delta_{b}^{a},\hskip 10.0pt$ $e_{\mu}{}^{a}e^{\nu}{}_{a}=\delta_{\mu}^{\nu}-\tau_{\mu}v^{\nu}.$ (33)
@@
$g_{\mu\nu}=\tfrac{-1}{\omega}\tau_{\mu}\tau_{\nu}+e_{\mu}{}^{a}e_{\nu}{}^{b}\delta_{ab},\hskip 10.0pt$ $g^{\mu\nu}=-\omega v^{\mu}v^{\nu}+e^{\mu}{}_{a}e^{\nu}{}_{b}\delta^{ab}.$ (34)
@@ III.4 Equivalence of PNR and GAG approaches
$g_{\mu\nu}=-\omega\tau_{\mu}\tau_{\nu}+e_{\mu}{}^{a}e_{\nu}{}^{b}\delta_{ab},\hskip 10.0pt$ $g^{\mu\nu}=-\tfrac{1}{\omega}v^{\mu}v^{\nu}+e^{\mu}{}_{a}e^{\nu}{}_{b}\delta^{ab}.$ (38)
This is the same form as in the PNR parameterization with $\omega=c^{2}$ We can also define a Galilei compatible covariant derivative as
@@
where $C^{\rho}_{\mu\nu}$ is a Galilei compatible connection, $\omega_{\mu}{}^{a}$, $\omega_{\mu}^{ab}$ act as spin connections, and $\partial_{\mu}$ is the partial derivative with respect to a generic coordinate system. Defining ${h_{\mu\nu}=e_{\mu}{}^{a}e_{\nu}{}^{b}\delta_{ab}}$, we can see that the simplest connection that satisfies the above conditions is
@@ IV.1 $f(R)$ gravity
where $R_{1},R_{2},R_{3}$ are the expansion terms shown explicitly in (3.2) in [64]; they are also displayed in (88). (The use of the connection does not affect the from of the expansion as shown in Sec. III, different connections will result in different forms for $R_{1}$,$R_{2}$,$R_{3}$).
@@
Here, we need to be extra careful since $j$ can be unbounded regardless of the value of $i$, and appears in the power of $c$. For a theory to have consistent Galilean and Carrollian limits, the quantity ${2k-2n+j}$ must be bounded from above and below.
@@ IV.1.1 Carrollian limit
The term where $i=1$ and $j=0$ have $l=-2$, $m=0$ and $n=1$, while the term where $i=2$ and $j=N$ has $l=N-4$, $m=0$ and $n=2$. Putting all together we get
@@
In order for the second term to couple to some higher order of the first term, we must set $N\geq 2$ otherwise, it will be pure quadratic gravity, i.e., not a modification of GR. This is the same condition derived in [85]. The magnetic limit can be derived from (48) by noting that the first term has $i=1$, $j=0$, $l=-2$. Solving $2n+m+\frac{l+2-j}{2}=i$, we get $m=1$ and $n=0$. The second term has $i=2$, $j=N$ and $l=N-4$. Again, by solving $2n+m+\tfrac{l+2-j}{2}=i$, we get $m=n=1$. Putting all together we obtain
@@
Here the minimum value of $2k-2n$ is $-2$ i.e, $l=-2$ (by choosing $k=0,n=1$), and since $12n+m+\frac{l}{2}=$ 1 therefore $m=0$, substituting in the formula we get
@@
by solving $12n+m=$ 1, with in this case $n$ must be zero and $m=1$. Thus,
@@ IV.1.2 Galilean limit
The first terms in the NLO has $i=1$, $j=0$, $l=2$. Solving $12n+m+\tfrac{l-j}{2}=i+$ 1, we get $n=0$, $m=1$. The second term has $i=2$, $j=N$ and $l=N+4$. Solving $12n+m+\tfrac{l-j}{2}=i+$ 1, we get $n=0$, $m=1$. Putting all together we get
@@
Here we have to solve $22n+m+1=$ 2. The solution is $n=0$ and $m=1$, and we get
@@
The $n-$th order in the Galilean expansion can be derived from the same maximization problem by replacing $l$ with $l-2n$.
@@ IV.2.2 Carrollian limit
which is again solvable by the simplex method given that $a,b$ are known. The solution is $z=N-2a-4b-2$. Thus, for the term to be a modification of GR, the condition on $N$ must be $N\geq 2a+4b$. If such term exists in the Lagrangian then the Lagrangian is a modification of the Carrollian limit of GR. The $n$-th order is given by the modified optimization problem with $z\to z+2n$.
We can see that the conditions on $N$ in the two limits are mutually exclusive. Thus, no such term (of the form (67)) with finite $a$, $b$ can be a GR modification in the two limits simultaneously. This implies that no Lagrangian of the form $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$ with polynomial $f$ can be a modification of GR in both limits simultaneously. In both limits we can recover the results from the previous section if we set =¯$\mathcal{R}=\bar{\mathcal{R}}=R$. We can recover quadratic gravity examples in the previous section by setting $a=2,b=0$ or $a=0,b=2$ or $a=b=1$.
@@ IV.3 $f(g_{\mu\nu},R_{\mu\nu\sigma\rho},\nabla_{\mu})$ gravity
¯$\bar{\mathcal{R}}^{c_{ij}}$ $=\sum_{\mathclap{\begin{subarray}{c}\sum\limits_{m=0}^{4}\bar{k}_{ijm}=c_{ij}\\ \bar{k}_{ijm}\geq 0\end{subarray}}}\frac{c_{ij}!}{\bar{k}_{ij1}!...\bar{k}_{ijm}!}\prod_{\bar{t}=1}^{4}c^{2\bar{k}_{ij\bar{t}}(\bar{t}-3)}(\bar{\mathcal{R}}_{\bar{t}})^{\bar{k}_{ij\bar{t}}},$
@@
×$\mathrel{\phantom{=}}\times\nabla_{1}^{\alpha_{ij}}\nabla_{3}^{\gamma_{ij}}\prod_{t=1}^{4}\prod_{\bar{t}=1}^{4}c^{2k_{it}(t-2)+2\bar{k}_{ij\bar{t}}(\bar{t}-3)-2\alpha_{ij}+2\gamma_{ij}+N}(\mathcal{R}_{t})^{k_{it}}\nabla_{2}^{\beta_{ij}}(\mathcal{R}_{\bar{t}})^{\bar{k}_{ij\bar{t}}}.$
@@ Appendix A List of formulae
[…]
@@
    $\mathcal{R}_{3}$ $=T_{\sigma}\mathcal{K}_{\mu\alpha}T_{(\nu}B_{\lambda)}{}^{\alpha}-T_{\sigma}\mathcal{K}_{\nu\alpha}T_{(\mu}B_{\lambda)}{}^{\alpha}+\nabla_{\nu}(T_{(\mu}B_{\nu)\sigma})-\nabla_{\mu}(T_{(\nu}B_{\nu)\sigma})$
@@
[…]
@@
    ¯$\bar{\mathcal{R}}_{1}$ $=-V^{\alpha}V^{\beta}V^{\mu}\Pi^{\gamma\nu}\nabla_{\mu}\mathcal{K}_{\nu}^{\rho}-2V^{\alpha}V^{\beta}V^{\mu}\Pi^{\gamma\nu}C^{\sigma}_{[\mu\nu]}\mathcal{K}_{\sigma}^{\rho}-\tfrac{1}{2}V^{\alpha}V^{\beta}V^{\rho}V^{\lambda}\mathcal{K}_{\sigma}^{\gamma}B_{\lambda}{}^{\sigma}-V^{\alpha}V^{\gamma}V^{\nu}\Pi^{\mu\beta}\nabla_{\nu}\mathcal{K}_{\mu}^{\rho}$ (85)
@@
    ¯$\bar{\mathcal{R}}_{2}$ $=V^{\alpha}V^{\beta}V^{\gamma}V^{\nu}B^{\sigma\rho}B_{\sigma\nu}+\tfrac{1}{2}V^{\alpha}V^{\beta}V^{\mu}\Pi^{\nu\gamma}\nabla_{\nu}(T_{\lambda}B_{\mu}{}^{\rho})-V^{\alpha}V^{\beta}V^{\mu}V^{\lambda}\Pi^{\nu\gamma}\nabla_{\mu}(T_{(\nu}B_{\lambda)}{}^{\rho})$
    $\mathrel{\phantom{=}}-V^{\alpha}V^{\beta}V^{\mu}\Pi^{\nu\gamma}B_{\sigma}{}^{\rho}C^{\sigma}_{[\nu\mu]}+\tfrac{1}{2}V^{\alpha}V^{\beta}B_{\sigma}{}^{\rho}\mathcal{K}^{\gamma\sigma}+V^{\alpha}V^{\gamma}V^{\lambda}V^{\nu}\Pi^{\mu\beta}\nabla_{\nu}(T_{(\mu}B_{\lambda)}{}^{\rho})$
@@
    $\mathrel{\phantom{=}}-V^{\beta}V^{\mu}\Pi^{\nu\gamma}\Pi^{\alpha\lambda}\nabla_{\nu}(T_{(\mu}B_{\lambda)}{}^{\rho})+V^{\beta}V^{\mu}\Pi^{\gamma\nu}\Pi^{\alpha\lambda}\nabla_{\mu}(T_{(\nu}B_{\lambda)}{}^{\rho})-V^{\gamma}V^{\nu}\Pi^{\mu\beta}\Pi^{\alpha\lambda}\nabla_{\nu}(T_{(\mu}B_{\lambda)}{}^{\rho})$
@@
    ¯$\bar{\mathcal{R}}_{4}$ $=\Pi^{\alpha\lambda}\Pi^{\mu\beta}\Pi^{\nu\gamma}\nabla_{\nu}(T_{(\mu}B_{\lambda)}{}^{\rho})-\Pi^{\alpha\lambda}\Pi^{\mu\beta}\Pi^{\nu\gamma}\nabla_{\mu}(T_{(\nu}B_{\lambda)}{}^{\rho})$
@@
[…]
@@
    ¯$\bar{\mathcal{R}}_{1}$ $=V^{\sigma}V^{\rho}V^{\nu}\nabla_{\nu}\mathcal{K}-2V^{\sigma}V^{\rho}V^{\nu}C^{\alpha}_{[\nu\beta]}\mathcal{K}_{\alpha}^{\beta}$ (87)
    ¯$\bar{\mathcal{R}}_{2}$ $=-V^{\sigma}V^{\rho}V^{\mu}\nabla_{\alpha}B_{\mu}{}^{\alpha}-V^{\rho}\Pi^{\mu\sigma}\nabla_{\alpha}\mathcal{K}_{\mu}^{\alpha}V^{\rho}V^{\nu}\Pi^{\nu\sigma}\nabla_{\nu}(T_{\mu}\mathcal{K})-V^{\rho}V^{\nu}\mathcal{K}^{\alpha\sigma}B_{\nu\alpha}+V^{\sigma}\Pi^{\nu\rho}\nabla_{\nu}\mathcal{K}$

### v4

@@ I Introduction
While calculating the Carrollian limit of Maxwell’s theory, it was shown that there are two nonequivalent Carrollian limits for any Lorentzian theory. The first one keeps the electric part in Maxwell’s theory and so was named the “electric limit”, and the other keeps the magnetic part and so was called the “magnetic limit”. In gravitational theories, the electric limit is the LO in the Carrollian expansion of the Lagrangian and the magnetic limit is the NLO in that expansion. There are multiple ways to take the (electric and magnetic) Carrollian limit of a Lorentzian gravitational theory:
@@
In this paper, we show that the PNR parametrization can be constructed from GAG. Similarly, the PUL parametrization can be constructed from CAG. We also show that the PUL and ZS approaches of taking the Carrollian limit are completely equivalent (and, similarly, PNR and IS). Throughout this paper, we consider the ADM decomposition with an extra real parameter $\epsilon$ which can be interpreted as a signature if it is $1$, 1-1, or 0; the zero signature case is a Carrollian manifold [79]. Once we establish the uniqueness of the Galilean and Carrollian limits (i.e., the equivalence of above approaches), we construct an algorithm for computing the $n$-th order of the Galilean, and the Carrollian expansions of a completely generic higher derivative gravity (HDG). The problem is transformed into computationally easier optimization problems which upon solving gives the desired order.
@@ II.4 Infinite scaling approach
Starting from a Lorentzian theory and writing the metric in the same form as in the ZS approach but taking the limit of the norm of the normal vector to -$\infty$ results in a Galilean theory. Alternatively, starting with the Riemannian theory and taking the limit to $\infty$ results in the same Galilean theory; the difference would be only in an overall sign in the Lagrangian.
@@ II.5 Carroll algebra gauging
$H\to\tau_{\mu},$ $P_{a}\to e_{\mu}{}^{a}$, $G_{a}\to\omega_{\mu}{}^{a}$, $\hskip 10.0ptJ_{ab}\to\omega_{\mu}{}^{ab}$. (11)
@@
$\zeta=\xi^{\mu}\tau_{\mu},$ $\zeta^{a}=\xi^{\mu}e_{\mu}{}^{a}$, $\hskip 10.0pt\lambda^{a},$ $\hskip 10.0pt\lambda^{ab}.$ (13)
@@
[…]
@@
[…]
@@ II.6 Galilei algebra gauging
[…]
@@ III.1 Equivalence of PUL and ZS approaches
This is the PUL form that appeared in [64] with $\epsilon=-c^{2}$. Thus, the limit $\epsilon\to 0$ gives the same result as the limit $c\to 0$. [^1] A Galilean analogue of this limit is to take $\epsilon\to$-$\infty$. This limit gives correct Galilean expansions. Given that the parameter $\epsilon$ is related to the norm of the normal vector. The Galilean limit corresponds to infinite normalization of this vector. This means that the two methods are equivalent. It was shown that the result would be a Carrollian manifold in [79].
@@
where $A_{1}^{\mu}=-n^{\mu}\partial_{\nu}n^{\nu}+n^{\nu}\partial_{\nu}n^{\mu}-\tfrac{1}{2}n^{\mu}n^{\sigma}h^{\alpha\rho}\big(2\partial_{(\alpha}h_{\sigma)\rho}-\partial_{\rho}h_{\alpha\sigma}\big)+h^{\mu\rho}n^{\sigma}n^{\nu}\partial_{\nu}h_{\sigma\rho}-\tfrac{1}{2}n^{\sigma}n^{\nu}h^{\mu\rho}\partial_{\rho}h_{\sigma\nu}$, and $A_{2}^{\mu}=-h^{\mu\rho}n^{\sigma}\partial_{\sigma}n_{\rho}$, where $\partial_{\mu}$ is the partial derivative with respect to a generic coordinate system. This formula is the same as the PUL parameterization of the Ricci scalar up to change of connections and renaming quantities.
@@ III.3 Equivalence of PUL and CAG approaches
$v^{\mu}\tau_{\mu}=-1,$ $\hskip 10.0ptv^{\mu}e_{\mu}{}^{a}=\tau_{\mu}e^{\mu}{}_{b}=0,\hskip 10.0pt$ $e_{\mu}{}^{a}e^{\mu}{}_{b}=\delta_{b}^{a},$ $e_{\mu}{}^{a}e^{\nu}{}_{a}=\delta_{\mu}^{\nu}-\tau_{\mu}v^{\nu}.$ (33)
@@
$g_{\mu\nu}=\tfrac{-1}{\omega}\tau_{\mu}\tau_{\nu}+e_{\mu}{}^{a}e_{\nu}{}^{b}\delta_{ab},$ $g^{\mu\nu}=-\omega v^{\mu}v^{\nu}+e^{\mu}{}_{a}e^{\nu}{}_{b}\delta^{ab}.$ (34)
@@ III.4 Equivalence of PNR and GAG approaches
$g_{\mu\nu}=-\omega\tau_{\mu}\tau_{\nu}+e_{\mu}{}^{a}e_{\nu}{}^{b}\delta_{ab},$ $g^{\mu\nu}=-\tfrac{1}{\omega}v^{\mu}v^{\nu}+e^{\mu}{}_{a}e^{\nu}{}_{b}\delta^{ab}.$ (38)
This is the same form as in the PNR parameterization with $\omega=c^{2}$. We can also define a Galilei compatible covariant derivative as
@@
where $C^{\rho}_{\mu\nu}$ is a Galilei compatible connection, $\omega_{\mu}{}^{a}$ and $\omega_{\mu}^{ab}$ act as spin connections, and $\partial_{\mu}$ is the partial derivative with respect to a generic coordinate system. Defining $h_{\mu\nu}=e_{\mu}{}^{a}e_{\nu}{}^{b}\delta_{ab}$}, we can see that the simplest connection that satisfies the above conditions is
@@ IV.1 $f(R)$ gravity
where $R_{1},R_{2},R_{3}$ are the expansion terms shown explicitly in (3.2) in [64]; they are also displayed in (88). (The use of the connection does not affect the from of the expansion as shown in Sec. III, different connections will result in different forms for $R_{1}$$R_{2}$$R_{3}$).
@@
Here, we need to be extra careful since $j$ can be unbounded regardless of the value of $i$, and appears in the power of $c$. For a theory to have consistent Galilean and Carrollian limits, the quantity 2k2n+j{2k-2n+j} must be bounded from above and below.
@@ IV.1.1 Carrollian limit
The term where $i=1$ and $j=0$ have l=2l=-2, $m=0$ and $n=1$, while the term where $i=2$ and $j=N$ has l=N4l=N-4, $m=0$ and $n=2$. Putting all together we get
@@
In order for the second term to couple to some higher order of the first term, we must set $N\geq 2$ otherwise, it will be pure quadratic gravity, i.e., not a modification of GR. This is the same condition derived in [85]. The magnetic limit can be derived from (48) by noting that the first term has $i=1$, $j=0$, l=2l=-2. Solving $2n+m+\frac{l+2-j}{2}=i$, we get $m=1$ and $n=0$. The second term has $i=2$, $j=N$ and l=N4l=N-4. Again, by solving $2n+m+\tfrac{l+2-j}{2}=i$, we get $m=n=1$. Putting all together we obtain
@@
Here the minimum value of 2k2n2k-2n is 2-2 i.e, l=2l=-2 (by choosing $k=0,n=1$), and since $2n+m+\frac{l}{2}=1$ therefore $m=0$, substituting in the formula we get
@@
by solving $2n+m=1$, with in this case $n$ must be zero and $m=1$. Thus,
@@ IV.1.2 Galilean limit
The first terms in the NLO has $i=1$, $j=0$, $l=2$. Solving $2n+m+\tfrac{l-j}{2}=i+1$, we get $n=0$, $m=1$. The second term has $i=2$, $j=N$ and $l=N+4$. Solving $2n+m+\tfrac{l-j}{2}=i+1$, we get $n=0$, $m=1$. Putting all together we get
@@
Here we have to solve $2n+m+1=2$. The solution is $n=0$ and $m=1$, and we get
@@
The $n$-th order in the Galilean expansion can be derived from the same maximization problem by replacing $l$ with l2nl-2n.
@@ IV.2.2 Carrollian limit
which is again solvable by the simplex method given that $a,b$ are known. The solution is z=N2a4b2z=N-2a-4b-2. Thus, for the term to be a modification of GR, the condition on $N$ must be $N\geq 2a+4b$. If such term exists in the Lagrangian then the Lagrangian is a modification of the Carrollian limit of GR. The $n$-th order is given by the modified optimization problem with $z\to z+2n$.
We can see that the conditions on $N$ in the two limits are mutually exclusive. Thus, no such term (of the form (67)) with finite $a$, $b$ can be a GR modification in the two limits simultaneously. This implies that no Lagrangian of the form $f(g_{\mu\nu},R_{\mu\nu\sigma\rho})$ with polynomial $f$ can be a modification of GR in both limits simultaneously. In both limits we can recover the results from the previous section if we set $\mathcal{R}=\bar{\mathcal{R}}=R$. We can recover quadratic gravity examples in the previous section by setting $a=2,b=0$ or $a=0,b=2$ or $a=b=1$.
@@ IV.3 $f(g_{\mu\nu},R_{\mu\nu\sigma\rho},\nabla_{\mu})$ gravity
$\bar{\mathcal{R}}^{c_{ij}}$ $=\sum_{\mathclap{\begin{subarray}{c}\sum\limits_{m=0}^{4}\bar{k}_{ijm}=c_{ij}\\ \bar{k}_{ijm}\geq 0\end{subarray}}}\frac{c_{ij}!}{\bar{k}_{ij1}!...\bar{k}_{ijm}!}\prod_{\bar{t}=1}^{4}c^{2\bar{k}_{ij\bar{t}}(\bar{t}-3)}(\bar{\mathcal{R}}_{\bar{t}})^{\bar{k}_{ij\bar{t}}},$
@@
$\mathrel{\phantom{=}}\times\nabla_{1}^{\alpha_{ij}}\nabla_{3}^{\gamma_{ij}}\prod_{t=1}^{4}\prod_{\bar{t}=1}^{4}c^{2k_{it}(t-2)+2\bar{k}_{ij\bar{t}}(\bar{t}-3)-2\alpha_{ij}+2\gamma_{ij}+N}(\mathcal{R}_{t})^{k_{it}}\nabla_{2}^{\beta_{ij}}(\mathcal{R}_{\bar{t}})^{\bar{k}_{ij\bar{t}}}.$
@@ Appendix A List of formulae
    $\mathcal{R}_{2}$ $=\overset{c}{R}_{\sigma\lambda\mu\nu}+2T_{\sigma}\nabla_{[\nu}\mathcal{K}_{\mu]\lambda}+\mathcal{K}_{\lambda\alpha}C^{\alpha}_{[\nu\mu]}T_{\sigma}+2T_{\sigma}T_{\lambda}\mathcal{K}^{\alpha}_{[\nu}\mathcal{K}_{\mu]\alpha}+2\nabla_{[\mu}(\mathcal{K}_{\nu]\sigma}T_{\lambda})$
@@
    $\mathcal{R}_{3}$ $=T_{\sigma}\mathcal{K}_{\mu\alpha}T_{(\nu}B_{\lambda)}{}^{\alpha}-T_{\sigma}\mathcal{K}_{\nu\alpha}T_{(\mu}B_{\lambda)}{}^{\alpha}+\nabla_{\nu}(T_{(\mu}B_{\lambda)\sigma})-\nabla_{\mu}(T_{(\nu}B_{\lambda)\sigma})$
@@
    where $B_{\mu\nu}=\partial_{\mu}T_{\nu}-\partial_{\nu}T_{\mu}$, $C^{\rho}_{\mu\nu}$ is the connection in (3), $\nabla_{\mu}$ is its compatible covariant derivative, and $\overset{c}{R}_{\sigma\lambda\mu\nu}$ is its Riemann tensor.
@@
    $\bar{\mathcal{R}}_{1}$ $=-V^{\alpha}V^{\beta}V^{\mu}\Pi^{\gamma\nu}\nabla_{\mu}\mathcal{K}_{\nu}^{\rho}-2V^{\alpha}V^{\beta}V^{\mu}\Pi^{\gamma\nu}C^{\sigma}_{[\mu\nu]}\mathcal{K}_{\sigma}^{\rho}-\tfrac{1}{2}V^{\alpha}V^{\beta}V^{\rho}V^{\lambda}\mathcal{K}_{\sigma}^{\gamma}B_{\lambda}{}^{\sigma}-V^{\alpha}V^{\gamma}V^{\nu}\Pi^{\mu\beta}\nabla_{\nu}\mathcal{K}_{\mu}^{\rho}$ (85)
@@
[…]
@@
    $\bar{\mathcal{R}}_{3}$ $=\overset{c}{R}{}^{\rho\alpha\beta\gamma}+\tfrac{1}{4}V^{\alpha}V^{\beta}B_{\sigma}{}^{\rho}B^{\sigma\rho}-V^{\alpha}V^{\lambda}\Pi^{\mu\beta}\Pi^{\nu\gamma}\nabla_{\nu}(T_{(\nu}B_{\lambda)}{}^{\rho})+V^{\alpha}V^{\lambda}\Pi^{\mu\beta}\Pi^{\nu\gamma}\nabla_{[\mu}\mathcal{K}_{\nu]}^{\rho}+\tfrac{1}{4}V^{\beta}V^{\gamma}B_{\sigma}{}^{\rho}B^{\alpha\sigma}$
@@
[…]
@@
    $\mathcal{R}_{2}$ $=\overset{c}{R}_{\mu\nu}+\nabla_{\sigma}(T_{\nu}\mathcal{K}_{\mu}^{\sigma})-\nabla_{\nu}(T_{\mu}\mathcal{K})+2C^{\rho}_{[\nu\sigma]}T_{\mu}\mathcal{K}_{\rho}^{\sigma}+\mathcal{K}_{(\mu}^{\alpha}B_{\nu)\alpha}-V^{\sigma}\mathcal{K}_{(\nu}^{\alpha}T_{\mu)}B_{\sigma\alpha}$
@@
    $\bar{\mathcal{R}}_{1}$ $=V^{\sigma}V^{\rho}V^{\nu}\nabla_{\nu}\mathcal{K}-2V^{\sigma}V^{\rho}V^{\nu}C^{\alpha}_{[\nu\beta]}\mathcal{K}_{\alpha}^{\beta}$ (87)
    $\bar{\mathcal{R}}_{2}$ $=-V^{\sigma}V^{\rho}V^{\mu}\nabla_{\alpha}B_{\mu}{}^{\alpha}-V^{\rho}\Pi^{\mu\sigma}\nabla_{\alpha}\mathcal{K}_{\mu}^{\alpha}V^{\rho}V^{\nu}\Pi^{\nu\sigma}\nabla_{\nu}(T_{\mu}\mathcal{K})-V^{\rho}V^{\nu}\mathcal{K}^{\alpha\sigma}B_{\nu\alpha}+V^{\sigma}\Pi^{\nu\rho}\nabla_{\nu}\mathcal{K}$